OpenStudy (anonymous):

Condense: ln40+2ln(1/2)+lnx

OpenStudy (anonymous):

ln means all + becomes * and - is /. so: 40*.5^2+x... .25*40=1+x.. x=-1

OpenStudy (anonymous):

Isn't it supposed to be an equation?

OpenStudy (anonymous):

40*(.5)^2+x?